A boat is travelling at 15 miles per hour. Aman walks across the deck in a direction at a right angle to the direction of the boat at the rate of 6ft. Per seconds. What is the total distance travelled by the man in 1 second with respect to the earth?
1
Expert's answer
2020-10-15T10:50:48-0400

velocity of boat (vb) =15 mile per hour

=22 ft/s

velocity of man(vm)=6 ft/s

so

displacement of man with respect to boat in 1 sec(sm) ="v_mt"

"s_m=6\times1=6 ft"


displacement of boat(sb)="v_bt"

"s_b=22\times1=22ft"

so displacement of man with respect to earth will be


"s=\sqrt{s_m^2+s_b^2}\\s=\sqrt{6^2+22^2}=22.8 ft"


so displacement of man with respect to earth will be 22.8ft